# Exact Hull Reformulation for Quadratically Constrained Generalized Disjunctive Programs

### v1

@@ 2.3.2 Numerical Challenges and $\varepsilon$-approximation
Equation (4) outlines the general strategy for handling constraints in Hull Reformulation. In the case of linear constraints of the form $A\mathbf{x}\leq b$, instead of using the definition of closure of the perspective function, the reformulation of linear constraints can be expressed simply as $A\mathbf{v}\leq by$ [5], which preserves linearity and exactly analytically describes the convex hull of the disjunction.
However, nonlinear constraints present a numerical challenge since the closure of the perspective function is not defined analytically at $y=0$. To overcome this issue, $\varepsilon$-approximation method (where $\varepsilon$ is a small number, e.g. $10^{-4}$) proposed by [12] is typically employed:
@@
Although computationally practical, $\varepsilon$-approximation results in exact Hull Reformulation only when $\varepsilon\rightarrow 0$, and however small, any values $\varepsilon>0$ enlarge the feasible region of the continuous relaxation [12], potentially weakening solver effectiveness. Additionally, application of $\varepsilon$-approximation of the closure of the perspective function creates a more complex functional structure compared to the original disjunctive constraints, introducing numerical instabilities, further motivating the development of an exact reformulation approach presented later in this work.
2.3.3 Convex and Non-Convex Constraints
When dealing with convex GDPs, where each disjunctive constraint is convex, the continuous relaxation of Hull Reformulation creates a convex hull of disjuncts in a disjunction [11], which is the tightest possible convex relaxation for the original disjunctive structure. This property is the key theoretical motivation behind the Hull Reformulation and is what makes it particularly attractive.
However, the behavior of the closure of the perspective function at binary values, described above, and thus the validity of reformulation, is independent of convexity, making it a valid reformulation to apply to non-convex GDPs (Appendix A). Applying Hull Reformulation directly to non-convex GDPs results in non-convex MINLP problems that can be solved by global optimization solvers.
Figure 1 illustrates the feasible region of continuous relaxation obtained from applying Hull Reformulation to a non-convex GDP, providing intuitive insight into the relaxation’s structure and its enlargement when using approximations. However, it should be noted that, since in this case the relaxation is not convex, the solver might need to apply its own techniques to obtain convex relaxations during the solution process to compute lower bounds, meaning that the convex region explored by the solver can be larger compared to the non-convex one obtained after applying the Hull Reformulation.
Figure 1: Illustration of the effect of the $\varepsilon$-approximation on the feasible region of the continuous relaxation resulting from the Hull Reformulation applied to a non-convex GDP. The figure compares the exact feasible region with the enlarged region obtained using the $\varepsilon$-approximation, highlighting how the approximation leads to a relaxation that is larger than the exact one and weakens as $\varepsilon$ increases. The Big-M relaxation is not shown, as its feasible region would significantly exceed the plotted area, taking up most of the figure at this scale and axis limits.
@@ 2.3.4 Trade-offs and Practical Considerations
The Hull Reformulation significantly improves the quality of the continuous relaxation, leading to tighter lower bounds. However, this comes at the cost of increased model complexity due to the introduction of additional variables and constraints [7]. This trade-off between stronger relaxations and higher dimensionality makes it difficult to determine in general whether the hull reformulation will outperform the Big-M reformulation in terms of solution time, and the results vary on a case-by-case basis [15, 3].
@@ 2.4 Binary Multiplication Reformulation
One of the simplest and most straightforward methods to transform the GDP problem into a MINLP formulation is the binary multiplication approach. This method involves directly multiplying each constraint within a disjunct by the binary variable associated with the disjunct, which activates or deactivates the corresponding constraint. Binary multiplication reformulation of disjunctions can be expressed as:
$\begin{aligned} &y_{ik}h_{ik}(\mathbf{x})\leq 0\quad\forall\,k\in K,\,i\in D_{k},\\ \end{aligned}\qquad\text{(Binary Multiplication Reformulation)}$ (7)
In this formulation, when the binary variable $y_{ik}=0$, the corresponding constraint becomes $0\leq 0$, which is trivially satisfied. When $y_{ik}=1$, the original constraint $h_{ik}(\mathbf{x})\leq 0$ is recovered. Therefore, the disjunctive logic of the original GDP formulation is correctly enforced, and the reformulation is valid.
However, the main drawback of this approach is that the continuous relaxation of the reformulation is often non-convex. In fact, the continuous relaxation has the same feasible region as the original MINLP with binary values, when projected to $\mathbf{x}$, providing no additional convex relaxation benefit to the solver. As a result, MINLP solvers must rely on internal convexification techniques to navigate the non-convex feasible space efficiently during the branch-and-bound procedure.
Despite its lack of favorable relaxation properties, the binary multiplication reformulation remains valid and easy to implement. It can be directly applied to GDP problems and passed to a general MINLP solver, which can handle non-convex problems.
3 Quadratically constrained GDP
In this work, we focus on the hull reformulation of GDPs with disjunctive quadratic constraints. Disjunctions of quadratically constrained GDP can be expressed as follows:
@@ 3 Quadratically constrained GDP
Application of Hull Reformulation to the quadratically constrained disjunctions, using $\varepsilon$-approximation of the closure of the perspective function shown in Equation (6), after algebraic simplification yields:
@@
4 Proposed Hull Reformulation
Consider a GDP in which the disjunctions contain quadratic constraints of the form $h(\mathbf{x})\;=\;\mathbf{x}^{\mathsf{T}}Q\mathbf{x}+\mathbf{c}^{\mathsf{T}}\mathbf{x}+d\leq 0,$ with continuous relaxation obtained through the standard exact Hull Reformulation. In lifted space, the feasible set of the continuous relaxation associated with a single constraint is
@@ 4 Proposed Hull Reformulation
Multiplying the function $\dfrac{\mathbf{v}^{\mathsf{T}}Q\mathbf{v}}{y}+c^{\mathsf{T}}\mathbf{v}+dy$ by $y$ motivates the alternative set:
@@
Case $y=0$: The constraints $\mathbf{x}^{\ell}y\leq\mathbf{v}\leq\mathbf{x}^{u}y$ force $\mathbf{v}=\mathbf{0}$ at $y=0$. Substituting $(\mathbf{0},0)$ into the $\mathbf{v}^{\top}Q\mathbf{v}+(\mathbf{c}^{\top}\mathbf{v})y+dy^{2}$ yields 0, identical to $\operatorname{cl}\widetilde{h}(\mathbf{0},0)$. Thus, the feasible points with $y=0$ are the same in both sets.
@@
Hence, the Hull Reformulation for quadratically constrained GDP problems can be exactly expressed using set $S_{2}$ for each quadratic constraint, avoiding the need for approximations and maintaining quadratic functional structure in the reformulation. Consequently, when applying Hull Reformulation it is proposed to use $S_{2}$ for any disjunctive quadratic constraints. In this case, the exact Hull Reformulation of quadratically constrained disjunctions shown in Equation (8) can be expressed as:
@@
The proposed exact reformulation provides two major advantages over $\varepsilon$-approximation approaches. First, since it is exact, it yields a tighter continuous relaxation, improving solver performance by delivering stronger lower bounds. Additionally, it preserves the quadratic structure of the problem, thereby enhancing computational efficiency and eliminating numerical instabilities associated with the $\varepsilon$-approximation. This structural preservation enables the use of advancements in Mixed-Integer Quadratically Constrained Programming (MIQCP) solvers when the original problem is quadratically constrained, which is not the case when $\varepsilon$-approximation is used, which alters the problem structure. Thus, we believe that the exact Hull Reformulation presented herein is preferable for any GDP problems involving quadratic constraints.
The proposed reformulation generalizes earlier results for specific quadratic constraints, such as $y\geq x^{2}$ [18], and for second-order cone–representable quadratic functions [13], applicable to convex quadratic function, and extends the approach to arbitrary quadratic functions, including non-convex ones.
Additionally, we note that this approach can be further generalized to polynomial constraints of arbitrary degree, enabling the construction of a Hull Reformulation in which the resulting expressions retain the same polynomial degree (Appendix B). Alternatively, polynomial constraints can be reformulated into an equivalent problem with quadratic constraints using established techniques [19], to which the proposed exact Hull Reformulation for quadratic constraints can be applied. Thus, using either strategy, the proposed exact Hull Reformulation approach can be extended to polynomial constraints of any degree.
5 Convex Quadratically Constrained GDP
@@
If $Q\succeq 0$ then $S_{1}$ is convex and therefore $S_{2}$ is also convex set by Proposition 1, even though $\mathbf{v}^{\top}Qv+(c^{\top}\mathbf{v})y+d\,y^{2}$ may not be a convex function in $(\mathbf{v},y)$. Some solvers detect convexity via the eigenvalue analysis of each quadratic expression and thus they may misclassify $S_{2}$ as non-convex, depending on the values of $\mathbf{c}$ and $d$.
If for convex quadratically constrained GDP it is desired to obtain a reformulation, that contains only convex functions, easily verified by eigenvalue analysis, it is possible to apply known convexification techniques to do so. Consider the following alternative convex set $S_{3}$, derived from set $S_{2}$ by relaxing the bilinear term $\mathbf{v}y$ through Glover’s linearization [20] and by substitution of $dy^{2}$ with $dy$, which is equivalent at binary $y$, to avoid non-convex function for $d<0$:
$S_{3}=\Big{\{}(\mathbf{v},y,\mathbf{z})$ $\in\mathbb{R}^{n}\times\mathbb{R}\times\mathbb{R}^{n}\;\Big{|}\;y\in[0,1],\;\mathbf{v}^{\top}Q\mathbf{v}+\mathbf{c}^{\top}\mathbf{z}+dy\leq 0,\;\mathbf{z}\leq\mathbf{x}^{u}y,\;\mathbf{z}\geq\mathbf{x}^{\ell}y,\;$ (14)
$\mathbf{z}\leq\mathbf{v}-\mathbf{x}^{u}(1-y),\;\mathbf{z}\geq\mathbf{v}-\mathbf{x}^{\ell}(1-y),\;\mathbf{x}^{\ell}y\leq\mathbf{v}\leq\mathbf{x}^{u}y\Big{\}}$
The transformed formulation utilizing $S_{3}$ maintains exact equivalence at binary values of $y$, ensures convexity of the functions, and can be verified by convexity detection methods. However, this explicitly convex reformulation may compromise solution efficiency because of the increased number of variables and constraints, as well as a potentially weaker continuous relaxation.
The set $S_{3}$ is included primarily to demonstrate that explicitly convex formulations are achievable and to illustrate one possible method by which solvers might internally manage the non-convex constraint $\mathbf{v}^{\top}Q\mathbf{v}+(\mathbf{c}^{\top}\mathbf{v})y+dy^{2}\leq 0$, when $Q\succeq 0$.
However, this approach can be less efficient because it weakens the continuous relaxation and increases computational overhead. Ideally, convexity detection techniques used by solvers should be applied directly to the original constraints in the GDP formulation, and the knowledge that certain constraints in the reformulation or the entire problem are convex should be retained and passed to a solver. This could eliminate unnecessary convexification steps or redundant relaxations that a solver may perform, thereby improving its computational performance. Unfortunately, most current solvers lack the capability to flag individual constraints or entire models as convex in this way, highlighting an area for improvement in solver design.
We believe that if solvers could recognize or be informed of the convexity of constraints mentioned above, or at least accept information indicating that the entire MINLP is convex when all disjunctive constraints are convex, the computational efficiency of the proposed exact Hull Reformulation for convex GDPs could be additionally improved compared to the results presented in this study.

### v2

@@ 2.2.3 Numerical Challenges and $\varepsilon$-approximation
Equation (4) outlines the general strategy for handling constraints in HR. In the case of linear constraints of the form $A\mathbf{x}\leq b$, instead of using the definition of closure of the perspective function, the reformulation of linear constraints can be expressed simply as $A\mathbf{v}\leq by$ [1], which preserves linearity and exactly analytically describes the convex hull of the disjunction.
However, nonlinear constraints present a numerical challenge since the closure of the perspective function is not defined analytically at $y=0$. To overcome this issue, $\varepsilon$-approximation method (where $\varepsilon$ is a small number, e.g. $10^{-4}$) proposed by Furman et al. [8], is typically employed:
@@
Although computationally practical and exact at binary values of $y$, $\varepsilon$-approximation results in exact HR for relaxation only when $\varepsilon\rightarrow 0$, and however small, any values $\varepsilon>0$ enlarge the feasible region of the continuous relaxation [8], potentially weakening solver effectiveness. Additionally, application of $\varepsilon$-approximation of the closure of the perspective function creates a more complex functional structure compared to the original disjunctive constraints, introducing numerical instabilities and making convexity detection more difficult, further motivating the development of exact HR approaches presented later in this work.
Figure 1 illustrates the feasible region of continuous relaxation obtained from applying HR to convex and non-convex GDPs, providing some intuition about the relaxation’s feasible region and its enlargement when $\varepsilon$-approximation is used.
(a) Convex GDP
(b) Non-convex GDP
Figure 1: Illustration of the effect of the $\varepsilon$-approximation on the feasible region of the continuous relaxation resulting from HR applied to convex and non-convex GDP. The figure shows the exact feasible region with the enlarged region obtained using the $\varepsilon$-approximation, highlighting how the approximation leads to a relaxation that is larger than the exact one and weakens as $\varepsilon$ increases. The Big-M relaxation is not shown, as its feasible region would significantly exceed the plotted area, taking up most of the figure at this scale and axis limits.
@@ 2.2.4 Trade-offs and Practical Considerations
HR significantly improves the quality of the continuous relaxation, leading to tighter lower bounds. However, this comes at the cost of increased model complexity due to the introduction of additional variables and constraints [11]. This trade-off between stronger relaxations and higher dimensionality makes it difficult to determine in general whether HR will outperform the Big-M reformulation in terms of solution time, and the results vary on a case-by-case basis [18, 17].
@@ 2.3 Binary Multiplication Reformulation
One of the simplest and most straightforward methods to transform the GDP problem into a MINLP formulation is the binary multiplication approach. This method involves directly multiplying each constraint within a disjunct by the binary variable associated with the disjunct, which activates or deactivates the corresponding constraint. The Binary Multiplication reformulation of disjunctions can be expressed as:
$\begin{aligned} &y_{ik}h_{ik}(\mathbf{x})\leq\mathbf{0}\quad\forall\,k\in K,\,i\in D_{k},\\ \end{aligned}\qquad\text{(Binary Multiplication Reformulation)}$ (7)
In this formulation, when the binary variable $y_{ik}=0$, the corresponding constraint becomes $\mathbf{0}\leq\mathbf{0}$, which is trivially satisfied. When $y_{ik}=1$, the original constraint $h_{ik}(\mathbf{x})\leq\mathbf{0}$ is recovered. Therefore, the disjunctive logic of the original GDP formulation is correctly enforced, and the reformulation is valid.
However, the main drawback of this approach is that the continuous relaxation of the reformulation is often non-convex. In fact, when projected to $\mathbf{x}$, the continuous relaxation has the same feasible region as the original MINLP with binary values, providing no additional convex relaxation benefit to the solver. As a result, for almost any problem, MINLP solvers must rely on internal convexification techniques to navigate the non-convex feasible space efficiently during the branch-and-bound procedure.
Despite its lack of favorable relaxation properties, the binary multiplication reformulation is valid and easy to implement. It can be directly applied to GDP problems and passed to a general MINLP solver, which can handle non-convex problems.
3 Hull Reformulations of Quadratically Constrained GDP
In this work, we focus on HR of GDPs with disjunctive quadratic constraints. Disjunctions of quadratically constrained GDP can be expressed as follows:
@@ 3 Hull Reformulations of Quadratically Constrained GDP
The application of HR to the quadratically constrained disjunctions, using $\varepsilon$-approximation of the closure of the perspective function shown in Equation (6), after algebraic simplification yields:
@@
3.1 General Exact Hull Reformulation of Quadratically Constrained GDP
Consider a GDP in which the disjunctions contain quadratic, possibly non-convex, constraints of the form $h(\mathbf{x})\;=\;\mathbf{x}^{\mathsf{T}}Q\mathbf{x}+\mathbf{c}^{\mathsf{T}}\mathbf{x}+d\leq 0,$ with finite bounds $\mathbf{x}^{\ell}\leq\mathbf{x}\leq\mathbf{x}^{u}$, and continuous relaxation obtained through the standard exact HR. In lifted space, the feasible set of the continuous relaxation associated with a single constraint is
@@ 3.1 General Exact Hull Reformulation of Quadratically Constrained GDP
Multiplying the function $\dfrac{\mathbf{v}^{\mathsf{T}}Q\mathbf{v}}{y}+\mathbf{c}^{\mathsf{T}}\mathbf{v}+dy$ by $y$ motivates the alternative set:
@@
Case $y=0$: The constraints $\mathbf{x}^{\ell}y\leq\mathbf{v}\leq\mathbf{x}^{u}y$ (finite bounds) force $\mathbf{v}=\mathbf{0}$ at $y=0$.
Substituting $(\mathbf{0},0)$ into $\mathbf{v}^{\top}Q\mathbf{v}+(\mathbf{c}^{\top}\mathbf{v})y+dy^{2}$ yields 0, identical to $\operatorname{cl}\widetilde{h}(\mathbf{0},0)$. Thus, the feasible points with $y=0$ are the same in both sets.
@@
Hence, HR for quadratically constrained GDP problems can be exactly expressed using set $S_{2}$ for each quadratic constraint, avoiding the need for approximations and maintaining quadratic functional structure in the reformulation. Consequently, when applying HR, it is possible to use $S_{2}$ for any disjunctive quadratic constraint. In this case, the General Exact Hull Reformulation (GEHR) of quadratically constrained disjunctions shown in Equation (8) can be expressed as:
@@
The proposed exact reformulation provides two major advantages over $\varepsilon$-approximation approaches. First, because it is exact, it avoids the feasible-region enlargement introduced by the $\varepsilon$-approximation and can therefore strengthen the relaxation and the resulting lower bounds. Additionally, it preserves the quadratic structure of the problem, potentially enhancing computational efficiency and eliminating numerical instabilities associated with the $\varepsilon$-approximation. This structural preservation enables the use of advancements in Mixed-Integer Quadratically Constrained Programming (MIQCP) solvers when the original problem is quadratically constrained, whereas the $\varepsilon$-approximation alters the problem structure.
However, the key drawback of this approach arises when this reformulation is applied to convex quadratic constraints. Although the resulting feasible region of the continuous relaxation is identical to that of the standard exact HR and is therefore convex, the inequality
$\mathbf{v}^{\top}Q\mathbf{v}+(\mathbf{c}^{\top}\mathbf{v})\,y+d\,y^{2}\leq 0$
is not necessarily convex as a function of $(\mathbf{v},y)$. As a result, solvers may classify these constraints as non-convex, which can significantly degrade performance compared to formulations that preserve explicit convexity, including the $\varepsilon$-approximation. We address this issue in the next section by re-deriving CEHR for the convex quadratic case.
Additionally, we note that this approach can be further generalized to polynomial constraints of arbitrary degree, enabling the construction of a HR in which the resulting expressions retain the same polynomial degree (Appendix B). Alternatively, polynomial constraints can be reformulated into an equivalent problem with quadratic constraints using established techniques [16], to which the proposed GEHR for quadratic constraints can be applied. Thus, using either strategy, the proposed GEHR approach can be extended to polynomial constraints of any degree.
3.2 Conic Exact Hull Reformulation of Convex Quadratically Constrained GDP
@@
It has been shown that HR of GDPs with second-order cone constraints admits an exact second-order cone representation [2]. However, to the best of our knowledge, the literature does not provide an explicit, solver-friendly algebraic formulation for HR of convex quadratic disjunctive constraints that avoids the perspective division and preserves a conic-representable structure. Moreover, this conic reformulation has not been integrated into commonly used GDP modeling frameworks such as Pyomo.GDP [7]. Thus, this subsection focuses on GDP disjunctions whose active constraints are convex quadratic inequalities.
Consider the following disjunctive constraint:
[…]
As it was mentioned before, the standard HR can be produced by the following constraint
$\bigl(\operatorname{cl}\widetilde{h}\bigr)(\mathbf{v},y)\;=\;\begin{cases}\dfrac{\mathbf{v}^{\mathsf{T}}Q\mathbf{v}}{y}+\mathbf{c}^{\mathsf{T}}\mathbf{v}+dy,&y>0,\\[6.0pt] 0,&y=0,\end{cases}\qquad\leq 0,$ (15)
together with the standard disaggregation bounds $\mathbf{x}^{\ell}y\leq\mathbf{v}\leq\mathbf{x}^{u}y$.
For convex quadratics, however, one can derive an exact algebraic formulation that avoids division and preserves quadratic structure, and is representable with a conic constraint.
3.2.1 Conic Exact Hull (without factorization)
The closure of the perspective function constraint (15) can be written (for $y>0$) as
$\frac{\mathbf{v}^{\mathsf{T}}Q\mathbf{v}}{y}+\mathbf{c}^{\mathsf{T}}\mathbf{v}+d\,y\;\leq\;0.$ (16)
A convenient way to avoid the division is to separate the quadratic-over-$y$ term from the affine terms via an epigraph variable. Introduce a scalar $t$ and impose
$\frac{\mathbf{v}^{\mathsf{T}}Q\mathbf{v}}{y}\;\leq\;t,\qquad t+\mathbf{c}^{\mathsf{T}}\mathbf{v}+d\,y\;\leq\;0.$ (17)
Because $Q\succeq 0$ implies $\mathbf{v}^{\mathsf{T}}Q\mathbf{v}\geq 0$, we may additionally enforce $t\geq 0$.
For $y>0$, the first inequality in (17) is equivalent to
$\mathbf{v}^{\mathsf{T}}Q\mathbf{v}\;\leq\;t\,y,\qquad t\geq 0,\;y>0,$ (18)
which is a (rotated) conic inequality. The second inequality in (17) is linear:
$t+\mathbf{c}^{\mathsf{T}}\mathbf{v}+d\,y\;\leq\;0.$ (19)
At $y=0$, the disaggregation bounds enforce $\mathbf{v}=\mathbf{0}$, so (18) holds trivially. Moreover, (19) together with $t\geq 0$ forces $t=0$, matching the closure value $\bigl(\operatorname{cl}\widetilde{h}\bigr)(\mathbf{0},0)=0$. Therefore, the projection of the system (18)–(19) (with $t\geq 0$ and the standard disaggregation bounds) onto $(\mathbf{v},y)$ coincides exactly with the closure of the perspective function constraint (15). In particular, this yields an exact reformulation that avoids the $\varepsilon$-approximation and removes the fractional term.
The validity of this reformulation is formally shown in the following proposition:
Proposition 2 (Conic epigraph reformulation of HR).
Let $Q\succeq 0$ and define
$S_{1}=\Bigl\{(\mathbf{v},y)\in\mathbb{R}^{n}\times\mathbb{R}\ \Bigm|\ y\in[0,1],\ \bigl(\operatorname{cl}\widetilde{h}\bigr)(\mathbf{v},y)\leq 0,\ \mathbf{x}^{\ell}y\leq\mathbf{v}\leq\mathbf{x}^{u}y\Bigr\},$ (20)
[…]
$S_{2}^{\mathrm{conic}}=\Bigl\{(\mathbf{v},y)\in\mathbb{R}^{n}\times\mathbb{R}\;\Bigm|$ $y\in[0,1],\ \exists\,t\in\mathbb{R},\ \mathbf{v}^{\mathsf{T}}Q\mathbf{v}\leq t\,y,$ (21)
$t+\mathbf{c}^{\mathsf{T}}\mathbf{v}+d\,y\leq 0,\ t\geq 0,\ \mathbf{x}^{\ell}y\leq\mathbf{v}\leq\mathbf{x}^{u}y\Bigr\}.$
Then $S_{1}=S_{2}^{\mathrm{conic}}$.
Proof.
We prove this by mutual containment of the sets.
First, take $(\mathbf{v},y)\in S_{1}$. If $y=0$, the bounds $\mathbf{x}^{\ell}y\leq\mathbf{v}\leq\mathbf{x}^{u}y$ imply $\mathbf{v}=\mathbf{0}$. Then the choice $t=0$ satisfies $\mathbf{v}^{\mathsf{T}}Q\mathbf{v}=0\leq t\,y=0$ and $t+\mathbf{c}^{\mathsf{T}}\mathbf{v}+d\,y=0$. Moreover, for any feasible $t$ we must have $t+\mathbf{c}^{\mathsf{T}}\mathbf{0}+d\cdot 0\leq 0$ and $t\geq 0$, hence $t=0$ is forced. Therefore $(\mathbf{v},y)\in S_{2}^{\mathrm{conic}}$. If $y>0$, define $t:=\frac{\mathbf{v}^{\mathsf{T}}Q\mathbf{v}}{y}\geq 0$ (since $Q\succeq 0$). Then $\mathbf{v}^{\mathsf{T}}Q\mathbf{v}\leq t\,y$ holds at equality, and $\bigl(\operatorname{cl}\widetilde{h}\bigr)(\mathbf{v},y)\leq 0$ implies $t+\mathbf{c}^{\mathsf{T}}\mathbf{v}+d\,y\leq 0$. Hence $(\mathbf{v},y)\in S_{2}^{\mathrm{conic}}$.
Second, take $(\mathbf{v},y)\in S_{2}^{\mathrm{conic}}$ with some $t\geq 0$ satisfying (21). If $y=0$, the bounds imply $\mathbf{v}=\mathbf{0}$, and then $\bigl(\operatorname{cl}\widetilde{h}\bigr)(\mathbf{0},0)=0\leq 0$, so $(\mathbf{v},y)\in S_{1}$. If $y>0$, the inequality $\mathbf{v}^{\mathsf{T}}Q\mathbf{v}\leq t\,y$ implies $\frac{\mathbf{v}^{\mathsf{T}}Q\mathbf{v}}{y}\leq t$, and combining with $t+\mathbf{c}^{\mathsf{T}}\mathbf{v}+d\,y\leq 0$ yields $\frac{\mathbf{v}^{\mathsf{T}}Q\mathbf{v}}{y}+\mathbf{c}^{\mathsf{T}}\mathbf{v}+d\,y\leq 0$, which is exactly $\bigl(\operatorname{cl}\widetilde{h}\bigr)(\mathbf{v},y)\leq 0$ for $y>0$. Therefore $(\mathbf{v},y)\in S_{1}$.
Thus $S_{1}=S_{2}^{\mathrm{conic}}$. ∎
Putting indices back and applying this construction to every convex quadratic constraint in each disjunct, the Conic Exact Hull Reformulation (CEHR) of the quadratically constrained disjunctions (8) is:
$\begin{aligned} &\mathbf{x}=\sum_{i\in D_{k}}\mathbf{v}_{ik},&&k\in K,\\ &\mathbf{x}^{\ell}y_{ik}\leq\mathbf{v}_{ik}\leq\mathbf{x}^{u}y_{ik},&&k\in K,\;i\in D_{k},\\ &\mathbf{v}_{ik}^{\mathsf{T}}Q_{jik}\mathbf{v}_{ik}\leq t_{jik}\,y_{ik},&&j\in J_{ik},\;k\in K,\;i\in D_{k},\\ &t_{jik}+\mathbf{c}_{jik}^{\mathsf{T}}\mathbf{v}_{ik}+d_{jik}\,y_{ik}\leq 0,&&j\in J_{ik},\;k\in K,\;i\in D_{k},\\ &t_{jik}\geq 0,&&j\in J_{ik},\;k\in K,\;i\in D_{k},\end{aligned}\qquad\text{(CEHR)}$ (22)
This formulation is exact for convex quadratic disjunctive constraints ($Q_{jik}\succeq 0$) and avoids the division by $y_{ik}$ present in the perspective form (15). Moreover, the only nonlinearities appear in the quadratic-over-product constraints $\mathbf{v}_{ik}^{\mathsf{T}}Q_{jik}\mathbf{v}_{ik}\leq t_{jik}y_{ik}$, which are rotated second-order cone representable. As a result, the HR can be handled by conic-capable solvers without explicitly factoring $Q_{jik}$ at the modeling layer.
3.2.2 Explicit Rotated Second Order Cone form via factorization
In our main implementation and experiments, we keep CEHR in the algebraic form shown in (22). Modern conic-capable solvers (notably Gurobi and SCIP) can often recognize a rotated second-order cone (RSOC) structure in (18) during presolve, so an explicit factorization of $Q$ is not required according to our observations. Additional information on the performance of different algebraic representations of conic constraints is provided in Appendix C.
Although we do not use the explicit conic form in our main computational study, we present it here for completeness. Since $Q\succeq 0$, there exists a matrix $L$ such that
$Q=LL^{\mathsf{T}}.$
For example, when $Q\succ 0$, $L$ can be obtained via a Cholesky factorization. When $Q$ is positive semidefinite (possibly singular), one option is a spectral factorization using an eigen-decomposition $Q=U\Lambda U^{\mathsf{T}}$ with $\Lambda\succeq 0$, yielding $L:=U\Lambda^{1/2}$.
Then
$\mathbf{v}^{\mathsf{T}}Q\mathbf{v}=\mathbf{v}^{\mathsf{T}}LL^{\mathsf{T}}\mathbf{v}=\|L^{\mathsf{T}}\mathbf{v}\|_{2}^{2},$
and (18) is equivalent to
$\|L^{\mathsf{T}}\mathbf{v}\|_{2}^{2}\leq t\,y,\qquad t\geq 0,\;y\geq 0.$
This is a rotated second-order cone representable set via
$\left(\frac{t}{2},\,y,\,L^{\mathsf{T}}\mathbf{v}\right)\in\mathcal{Q}_{r}\;:=\;\left\{(\alpha,\beta,\mathbf{w})\ \middle|\ 2\alpha\beta\geq\|\mathbf{w}\|_{2}^{2},\ \alpha\geq 0,\ \beta\geq 0\right\}.$
This yields an explicit mixed-integer second-order cone (MISOCP) formulation of HR for convex quadratic disjunctive constraints, re-deriving the known result that hull reformulation of second-order cone constraints admit second-order cone representations [2].
3.2.3 Extension to general second-order cone representable constraints
It has been shown in the literature that HR remains representable in the same cone families as those used within the original disjunctions [2]. Our convex quadratic case fits naturally within this framework. Although a generic convex quadratic inequality
$\mathbf{x}^{\mathsf{T}}Q\mathbf{x}+\mathbf{c}^{\mathsf{T}}\mathbf{x}+d\leq 0,\qquad Q\succeq 0,$
is not itself written as a cone membership constraint; it is second-order cone representable after introducing lifting.
In this paper, we focus on the conic reformulation of convex quadratic disjunctive constraints. However, the similar reformulation principle extends to any disjunctive constraint that admits an SOC representation, either in the original variables or after lifting. Thus, while our computational study for CEHR is restricted to convex quadratic constraints, the approach applies more broadly to SOC-representable disjunctive constraints, possibly after introducing auxiliary variables. The presented CEHR for convex functions can therefore be viewed as a special case of a more general result for conic-representable disjunctive constraints.